OpenStudy (anonymous):

What is one of the solutions to the following system of equations? y^2+x^2=53 y-x=5 a)(-10,-5) b)(-7,-2) c)(7,2) d)(5,10)

OpenStudy (solomonzelman):

based on the second equation, \(\large\color{slate}{y=x+5 }\) So, substitute this into the first equation. \(\large\color{slate}{(x+5)^2+x^2=53 }\) \(\large\color{slate}{x^2+10x+25+x^2=53 }\) \(\large\color{slate}{2x^2+10x-28 =0 }\) \(\large\color{slate}{x^2+5x-7 =0 }\) and then you finish...

OpenStudy (anonymous):

thank you